OpenStudy (anonymous):

show that 4x^2+9y^2=36 represent an ellipse find its directrix and eccentricity

OpenStudy (anonymous):

ok so i show that it is an ellipse by dividing throughout by 36 and simplifying to get \[\frac{ x^2 }{ 9 }+\frac{ y^2 }{ 4 }=1\]

OpenStudy (anonymous):

how do i find the directrix and eccentricity?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Hey!

ganeshie8 (ganeshie8):

From the standard form, can you figure out the values of \(a\) and \(b\) ?

ganeshie8 (ganeshie8):

Once you have \(a\) and \(b\), you may use below definition to find the value of eccentricity. http://www.mathopenref.com/ellipseeccentricity.html

Directrix (directrix):

An ellipse does not have a directrix.

OpenStudy (anonymous):

yea it does ..well that what was ask for the question

OpenStudy (anonymous):

c^2=9-4 c=sq.rt 5 so the focus is (sq.rt5,0) for the eccentricity i got c/a which gives (sq.rt 5)/3 for the directrix i got x=a^2/c which gives 9/sq.rt 5 is this correct

OpenStudy (anonymous):

@ganeshie8

OpenStudy (phi):

yes, that looks good

OpenStudy (anonymous):

ok thanks